# Shuffle-R1: Efficient RL framework for Multimodal Large Language Models via Data-centric Dynamic Shuffle

### v3

@@ Abstract
Reinforcement learning (RL) has emerged as an effective post-training paradigm for enhancing the reasoning capabilities of multimodal large language model (MLLM). However, current RL pipelines often suffer from training inefficiencies caused by two underexplored issues: Advantage Collapsing, where most advantages in a batch concentrate near zero, and Rollout Silencing, where the proportion of rollouts contributing non-zero gradients diminishes over time. These issues lead to suboptimal gradient updates and hinder long-term learning efficiency. To address these issues, we propose Shuffle-R1, a simple yet principled framework that improves RL fine-tuning efficiency by dynamically restructuring trajectory sampling and batch composition. It introduces (1) Pairwise Trajectory Sampling, which selects high-contrast trajectories with large advantages to improve gradient signal quality, and (2) Advantage-based Batch Shuffle, which increases exposure of valuable rollouts through strategic batch reshuffling. Experiments across multiple reasoning benchmarks demonstrate that our framework consistently outperforms strong RL baselines with minimal computational overhead. These results highlight the importance of data-centric adaptations for more efficient RL training in MLLM.
footnotetext: 🖂 Corresponding author.
@@ 1 Introduction
Inspired by the slow and deliberate thinking process in human cognition, reasoning capabilities empower Large Language Model (LLM) to plan, reflect, and generalize beyond mere memorization (Chu et al. 2025a; Kang et al. 2024). Recent advances have demonstrated that integrating reinforcement learning (RL) into the training of LLMs can significantly enhance such capabilities, particularly in complex domains such as mathematical problem solving and code generation (Guo et al. 2025; Deepmind 2025; OpenAI 2024; Seed et al. 2025). Notably, DeepSeek-R1 (Guo et al. 2025) leverage reward signals derived exclusively from verifiable outcomes to yield impressive performance gains. Beyond textual tasks, RL has also seen increasing application in multimodal domains, including object detection, segmentation, and video understanding (Li et al. 2025; Liu et al. 2025c, b; Wang et al. 2025c), highlighting its potential to support generalizable reasoning across modalities.
Figure 1: (a): Advantage Collapsing, where most advantages concentrate near zero. (b): Rollout Silencing, where the ratio of rollouts with non-zero gradient consistently drops. The phenomenon gets worse in larger models.
Despite the growing interest in RL for LLM and Multimodal LLM (MLLM), several practical challenges remain unresolved, including slow convergence, training instability, and suboptimal efficiency. Recent studies have proposed various improvements, such as empirical optimization of training configurations (Yu et al. 2025), integration of experience replay (Wang et al. 2025a), and refinement of objective functions (Chu et al. 2025b). Other works focus on reward optimization (Liu et al. 2025a; Ma et al. 2025).
However, most of these approaches remain confined to the standard RL training paradigm, where trajectories are uniformly sampled without considering their varying informativeness or difficulty. This static strategy overlooks a crucial insight that not all learning signals are equally valuable, which leads us to consider a fundamental question: Can dynamically prioritizing trajectories provide richer gradient information and lead to more efficient training?
@@
Motivated by this question, we examine the current RL training practices and identify two critical yet might underexplored inefficiencies. First, Advantage Collapsing emerges when most computed advantages cluster excessively near zero, drowning out informative signals from trajectories with large-magnitude advantages, resulting in weak or negligible gradient updates (Fig. 1(a)). Second, Rollout Silencing arises as the fraction of rollouts contributing non-zero gradients steadily declines during training (Fig. 1(b)), leading to wasted computation and underutilization of data. These two findings demonstrate a pressing need for adaptive mechanisms to prioritize, reuse, and reallocate gradient exposure toward informative samples.
@@
In this paper, we introduce Shuffle-R1, an innovative framework designed to dynamically prioritize and amplify critical gradient information. Rather than asking how to design better rewards, we try to ask what data should the model actually update on during RL fine-tuning. Specifically, Shuffle-R1 introduces two elegantly simple yet highly effective modules: (1) Pairwise Trajectory Sampling, which selects high-contrast trajectory pairs with large-magnitude advantages from an extended rollout pool, concentrating learning signals to mitigate Advantage Collapsing; and (2) Advantage-based Batch Shuffle, which dynamically reshapes training batches to prioritize informative trajectories while down-weighting ineffective ones, alleviating Rollout Silencing to improve data utilization. Our framework is lightweight, modular, and seamlessly integrable with existing RL algorithms. Experiments demonstrate that Shuffle-R1 significantly improves model performance across challenging multimodal reasoning tasks, even surpassing GPT-4o and Claude-3.7 on MathVerse and MathVista. Moreover, it achieves competitive performance to GRPO while requiring only half of the training steps.
In summary, our contributions are three-fold: 1) We reveal two critical yet underexplored limitations that undermine training efficiency in RL fine-tuning for MLLM, i.e., Advantage Collapsing and Rollout Silencing. 2) We propose Shuffle-R1, a novel and adaptive RL framework that dynamically selects high-contrast trajectories and reshapes training batches to emphasize informative samples. 3) Extensive experiments across model scales and both in-domain and out-of-domain benchmarks demonstrate the effectiveness and generalizability of our framework.
@@ 2.1 Large Reasoning Models
Researchers have explored various approaches to equip LLM with reasoning ability. Some early studies performed SFT on complex long chain-of-thought data, leading to performance gains on reasoning tasks (Muennighoff et al. 2025; Ye et al. 2025; Yue et al. 2024; Guo et al. 2024). However, some claim that SFT merely enables the model to memorize the format of reasoning steps and long chains of thought, without fully grasping the ability to reason independently (Chu et al. 2025a; Kang et al. 2024; Ye et al. 2024; Allen-Zhu and Li 2024). Some researchers control the model to generate structured chain-of-thought instead of free generation, achieving systematic step-by-step reasoning output (Xu et al. 2025; Wu et al. 2025; Thawakar et al. 2025) Other works attempted to use test-time scaling like Monte Carlo Tree Search (MCTS) (Yao et al. 2023; Zhang et al. 2024a; Du et al. 2024; Yao et al. 2024) to facilitate complex reasoning by actively extending the output of the model.
Recently, models such as OpenAI o1/o3 (OpenAI 2024), DeepSeek-R1 (Guo et al. 2025), Seed-Thinking (Seed et al. 2025), and Kimi-k1.5 (Team et al. 2025) utilized RL to enable the model to explore independently, stimulating reasoning ability. In particular, DeepSeek-R1-Zero directly conducted RL on pre-trained model without instruction fine-tuning, with verifiable outcome reward functions to replace reward models, achieving surprising reasoning ability. The training algorithms for RL are also constantly being optimized (Yu et al. 2025; Chu et al. 2025b). Our work focuses on a deeper investigation of the efficiency of RL training and puts forward an effective solution to improve both the efficiency and performance of RL training.
@@ 2.2 Reinforcement Learning for MLLM
Following the success of DeepSeek-R1, a series of studies have transplanted RL into the training of MLLM and downstream visual tasks, such as Open Vocabulary Object Detection (Liu et al. 2025c), Reasoning Segmentation (Liu et al. 2025b), Video Understanding (Li et al. 2025), Video Localization (Wang et al. 2025c), etc. These works mainly focus on the applicability of RL in downstream tasks. Some other works focusing on improving the general reasoning ability of MLLM have achieved performance improvement on reasoning tasks by collecting a large amount of high-quality data (Meng et al. 2025; Huang et al. 2025; Yang et al. 2025; Zhang et al. 2025; Peng et al. 2025; Tan et al. 2025). These works mainly focus on the organization of high-quality reasoning data and the balance between SFT and RL in the training process.
Some researchers who conduct in-depth research on the RL mechanism have optimized the RL training process from various aspects, including adding contrastive reward mechanism (Li et al. 2025), actively introducing reflection tokens during rollouts (Wang et al. 2025a), optimizing the RL objective function and gradient update mechanism (Chu et al. 2025b), introducing more diverse rollouts (Liu et al. 2025a; Yao et al. 2025), etc. The core objective of these works is to optimize the RL training process. In our work, we propose a novel training framework that introduces dynamic and adaptive selection and resampling to queries and rollouts, reshaping data distribution for better training efficiency and model performance.
3 Preliminaries
Reinforcement Learning in LLM.
Reinforcement learning is originally introduced as a post-training technique to align LLM outputs with human preference. Recently, it has been used in fine-tuning stage to improve the reasoning ability of LLMs. Policy gradient algorithm is the most commonly used RL method, whose objective function is to maximize the expectation of return the model gets from the environment. Given a query $q$, multiple independent sampled responses $O=\{o_{1},o_{2},...,o_{N}\}$ are sampled from the old policy model $o\sim\pi_{\theta^{\prime}}(q)$, with rollout size $N$. Subsequently, the corresponding rewards $R=\{r_{1},r_{2},...,r_{N}\}$ are obtained for each response, in our case they are calculated by verifiable reward functions. Advantages $\hat{A}=\{A^{\pi_{\theta}}(r_{1}),A^{\pi_{\theta}}(r_{2}),...,A^{\pi_{\theta}}(r_{N})\}$ are calculated to assess the improvement over the current policy model. Importance sampling and clipping are introduced to stabilize the training. The core objective function is defined as:
$\begin{split}&\mathcal{J}(\theta)=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{N}\sim\pi_{\theta^{\prime}}(\cdot|q)}\\ &\frac{1}{\sum_{i=1}^{N}|o_{i}|}\sum_{i=1}^{N}\sum_{t=1}^{|o_{i}|}\left\{\min\left[\gamma_{t}(\theta),\text{clip}\left(\gamma_{t}(\theta),1-\epsilon,1+\epsilon\right)\right]\hat{A}_{i}\right\}\end{split},$ (1)
@@ 3 Preliminaries
$\gamma_{t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta^{\prime}}(o_{i,t}|q,o_{i,<t})},$ (2)
$\hat{A_{i}}=\frac{r_{i}-\text{mean}(R)}{\text{std}(R)},$ (3)
and $\epsilon$ is clipping hyperparameter to prevent training collapse.
However, current static RL training paradigm still exhibits limitations. Most of the collected rollouts have advantages sharply concentrate near zero, providing limited signals for gradient update. The proportion of rollouts with non-zero gradient continues to decline throughout the training, introducing invalid updates. These limitations indicate the need for a more efficient training framework.
4 Method
In this section, we begin by further analyzing the existing inefficient issues in training. Then, as illustrated in Fig. 3, we introduce Shuffle-R1, which optimizes the training process through two crucial modules: (1) Pairwise Trajectory Sampling and (2) Advantage-based Batch Shuffle.
@@ 4.1 Problem Analysis
Our probe analysis reveals that, contrary to the ideal scenario, most rollouts exhibit advantages sharply concentrated around zero in current RL paradigm, leading to the Advantage Collapsing phenomenon. Such distribution results in weak gradient signals and hampers effective policy updates, as only rollouts with high-magnitude advantages can provide significant gradient signals. Simply increasing the number of rollouts can collect more amount of valuable rollouts and improve performance to some extent (Fig. 2(a)), but also substantially increase the computational overhead without fundamentally resolving the Advantage Collapsing. These findings highlight the need for a dynamic mechanism that can adaptively identify and select valuable rollouts to improve training efficiency.
@@
We observe that the proportion of rollouts contributing non-zero gradients steadily decreases as training progresses, resulting in the Rollout Silencing phenomenon. This is primarily caused by the accumulation of factors such as zero advantages, gradient clipping, and overlong truncation, which further exposes the limitations of static sampling paradigm. By tracking queries of varying difficulty, we find that simple queries tend to converge early, while difficult ones consistently yield low accuracy (Fig. 2(b)). Both cases fail to generate diverse and informative rollouts throughout training, contributing to Rollout Silencing. Moreover, in standard training pipelines, each rollout is used only once for a single gradient update, limiting the model to fully exploit valuable rollouts. Consequently, a dynamic sampling strategy that can adaptively discard ineffective rollouts and queries, while allowing repeated learning from valuable rollouts is of great importance.
Figure 2: (a) Model accuracy improves with larger rollout sizes. (b) Queries with different difficulties demonstrate varying accuracy during training, their corresponding rollouts have different diversity and qualities consequently.
@@ 4.2 Pairwise Trajectory Sampling
To relieve Advantage Collapsing, we seek to select trajectories that offer stronger learning signals. Since increasing the rollout number enlarges the chance of capturing high-advantage samples, we propose Pairwise Trajectory Sampling (PTS), a data-centric module to selectively amplify valuable learning signals. Rather than evaluating trajectories in isolation, PTS organizes candidate rollouts into structured contrastive pairs. This pairing mechanism captures both high and low advantage signals jointly, forming informative “positive-negative” pairs. Only pairs with the largest advantage contrast are then retained for training. This process ensures that limited update bandwidth is focused on trajectories that are both diverse and gradient-rich.
Given a query $q$ and a rollout size of $2N$, the rollout trajectories group is denoted as $O=\{o_{i}\}_{i=1}^{2N}$. The corresponding reward and advantage sets are $R=\{r_{i}\}_{i=1}^{2N}$ and $A=\{\hat{A}_{i}\}_{i=1}^{2N}$, respectively. To identify informative trajectory pairs, our proposed pairing mechanism follows a simple ‘max-and-min’ principle by matching the trajectory with the highest advantage to that with the lowest, the second highest to the second lowest, and so on. We denote the sorted advantage values in descending order as:
@@
In this scheme, the original $2N$ rollouts are easily sorted and reorganized into $N$ pairs. The top-ranked pairs typically consist of trajectories with high-magnitude but opposite-sign advantages, forming contrastive pairs akin to “positive-negative” samples. In contrast, the bottom-ranked pairs involve trajectories with advantages closer to zero.
@@
where $\alpha$ is a hyperparameter controlling the sampling ratio from the pairing set. We only keep the trajectories in the valid set for gradient update.
By introducing a structured contrastive sampling scheme, PTS enables more effective trajectory selection from a broader exploration space without increasing the gradient computation cost. The contrastive pairing not only filters out low-signal trajectories, but also sharpen the model’s policy gradient through direct comparison. PTS shifts the focus of RL fine-tuning from uniform exploration to gradient-informed selection, representing a principled step toward more efficient data usage in RL training.

### v4

@@ Abstract
Reinforcement learning (RL) has emerged as an effective post-training paradigm for enhancing the reasoning capabilities of multimodal large language model (MLLM). However, current RL pipelines often suffer from training inefficiencies caused by two underexplored issues: Advantage Collapsing, where most advantages in a batch concentrate near zero, and Rollout Silencing, where the proportion of rollouts contributing non-zero gradients diminishes over time. These issues lead to suboptimal gradient updates and hinder long-term learning efficiency. To address these issues, we propose Shuffle-R1, a simple yet principled framework that improves RL fine-tuning efficiency by dynamically restructuring trajectory sampling and batch composition. It introduces (1) Pairwise Trajectory Sampling, which selects high-contrast trajectories with large advantages to improve gradient signal quality, and (2) Advantage-based Batch Shuffle, which increases exposure of valuable rollouts through strategic batch reshuffling. Experiments across multiple reasoning benchmarks show that our framework consistently outperforms strong RL baselines with minimal computational overhead. These results support the potential of data-centric adaptations for more efficient RL training for MLLM.
@@ 1 Introduction
Reinforcement learning (RL) has emerged as a powerful tool to enhance large language models (LLMs) to plan, reflect, and generalize, enabling stronger performance in complex reasoning domains such as mathematical problem solving and code generation (Guo et al., 2025; Deepmind, 2025; OpenAI, 2024; Seed, 2025). Notably, DeepSeek-R1 (Guo et al., 2025) leverages reward signals derived exclusively from verifiable outcomes to yield impressive performance gains. Beyond textual tasks, RL has also seen increasing application in various multimodal domains (Li et al., 2025; Liu et al., 2025c; b; Wang et al., 2025c), highlighting its potential to support generalizable reasoning across modalities.
To better incorporate RL in LLM and Multimodal LLM (MLLM), recent studies have proposed various improvements, including better framework optimization (Yu et al., 2025; Wang et al., 2025a; Chu et al., 2025b) and more sophisticated reward optimization (Liu et al., 2025a; Ma et al., 2025). However, most approaches remain confined to static sampling paradigm, where trajectories are sampled and treated uniformly. Such strategy overlooks a crucial insight that not all learning signals are created equal and their informativeness varies and evolves during training. Ignoring variation in signal quality risks the training process being overwhelmed by noisy trajectories and underusing the truly useful signals. An ideal framework should instead concentrate updates on golden signals while filtering out noisy ones. Consequently, it leads us to consider a fundamental question: Can dynamically prioritizing trajectories provide richer gradient information and lead to more effective training?
@@
Motivated by this question, we investigate into current RL training practices and reveal two critical yet might underexplored limitations. First, Advantage Collapsing emerges when most computed advantages cluster excessively near zero, drowning out informative signals from trajectories with large-magnitude advantages, resulting in extremely weak or negligible gradient updates and wiping out decisive ones (Fig. 2(a)). Second, Rollout Silencing arises as the fraction of rollouts contributing non-zero gradients steadily declines during training (Fig. 2(b)), leading to catastrophic waste of computation without fully utilizing informative signals. These two findings demonstrate a pressing need for adaptive mechanisms to prioritize, reuse, and reallocate gradient exposure toward informative samples.
Figure 1: (a) Advantage Collapsing, where most advantages concentrate near zero. (b) Rollout Silencing, where the ratio of rollouts with non-zero gradient consistently drops.
Figure 2: Pipeline comparison. (a) Static paradigm. (b) Rule-based pre-filter paradigm. (c) Dynamic paradigm can ‘interact’ with model during training.
@@
In this paper, we present Shuffle-R1, a framework that dynamically prioritizes and amplifies critical gradient signals during RL fine-tuning. Guided by the philosophy that what data the model updates on is as important as how it updates, Shuffle-R1 introduces two effective modules: (1) Pairwise Trajectory Sampling, which selects high-contrast trajectory pairs with large advantages gaps from an extended rollout pool, concentrating on the most discriminative learning signals to mitigate Advantage Collapsing; and (2) Advantage-based Batch Shuffle, which adaptively reshapes training batches to emphasize informative trajectories while down-weighting unhelpful ones, alleviating Rollout Silencing by removing noisy signals and improving computation utilization. Together, our framework embodies the principle of dynamic data prioritization (Fig. 2), enabling adaptive interaction between model and data. Experiments show that Shuffle-R1 substantially improves multimodal reasoning performance, surpassing GPT-4o (Achiam et al., 2024) and Claude-3.7 (Anthropic, 2025) on MathVerse (Zhang et al., 2024b) and MathVista (Lu et al., 2023), and matches GRPO (Shao et al., 2024) with only half the training steps. In summary, our contributions are three-fold:
  • We reveal two critical yet underexplored limitations that undermine training efficiency in RL fine-tuning for MLLM, i.e., Advantage Collapsing and Rollout Silencing.
  • We propose Shuffle-R1, a novel and adaptive RL framework that dynamically selects high-contrast trajectories and reshapes training batches to emphasize informative samples.
  • Extensive experiments across model scales and both in-domain and out-of-domain benchmarks demonstrate the effectiveness and generalizability of our framework.
These results underscore the importance of rethinking which data to update on in RL post-training, moving beyond reward design toward dynamic and adaptive data structuring for more effective reasoning enhancement.
@@ 2.1 Large Reasoning Models
Researchers have explored various approaches to equip LLM with reasoning ability. Some early studies performed SFT on complex long chain-of-thought data, leading to performance gains on reasoning tasks (Muennighoff et al., 2025; Ye et al., 2025; Guo et al., 2024). However, it has been argued that SFT merely enables the model to memorize the format of reasoning steps and long chains of thought, without fully grasping the ability to reason independently (Chu et al., 2025a; Kang et al., 2024). Some researchers control the model to generate structured chain-of-thought instead of free generation, achieving systematic step-by-step reasoning output (Xu et al., 2025; Wu et al., 2025; Thawakar et al., 2025) Other works attempted to use test-time scaling like Monte Carlo Tree Search (MCTS) (Yao et al., 2023; Zhang et al., 2024a; Yao et al., 2024) to facilitate complex reasoning by actively extending the output of the model.
Recently, models such as OpenAI o1/o3 (OpenAI, 2024), DeepSeek-R1 (Guo et al., 2025), Seed-Thinking (Seed, 2025), and Kimi-k1.5 (Kimi, 2025) utilized RL to enable the model to explore independently, stimulating reasoning ability. In particular, DeepSeek-R1-Zero directly conducted RL on pre-trained model without instruction fine-tuning, with verifiable outcome reward functions to replace reward models, achieving remarkable reasoning ability. The training algorithms for RL are also constantly being optimized (Yu et al., 2025; Chu et al., 2025b). Our work focuses on a deeper investigation of the efficiency of RL training and proposes an effective solution to improve both the efficiency and performance of RL training.
@@ 2.2 Reinforcement Learning for MLLM
Following the success of DeepSeek-R1, a series of studies have transplanted RL into the training of MLLM and downstream visual tasks, such as Open Vocabulary Object Detection (Liu et al., 2025c), Reasoning Segmentation (Liu et al., 2025b), Video Understanding (Li et al., 2025), Video Localization (Wang et al., 2025c), etc. These works mainly focus on the applicability of RL to downstream tasks. Some other works focusing on improving the general reasoning ability of MLLM have achieved performance improvements on reasoning tasks by collecting a large amount of high-quality data (Meng et al., 2025; Huang et al., 2025; Yang et al., 2025; Zhang et al., 2025; Peng et al., 2025; Tan et al., 2025). These works mainly focus on the organization of high-quality reasoning data and the balance between SFT and RL in the training process.
Some researchers who conduct in-depth research on the RL mechanism have optimized the RL training process from various aspects, including adding contrastive reward mechanism (Li et al., 2025), actively introducing reflection tokens during rollouts (Wang et al., 2025a), optimizing the RL objective function and gradient update mechanisms (Chu et al., 2025b), and introducing more diverse rollouts (Liu et al., 2025a; Yao et al., 2025). The core objective of these works is to optimize the RL training process. In our work, we propose a novel training framework that introduces dynamic and adaptive selection and resampling of queries and rollouts, reshaping the data distribution for better training efficiency and model performance.
3 Method
In this section, we begin by further analyzing the existing drawbacks in training. Then, as illustrated in Fig. 4, we introduce Shuffle-R1, which optimizes the training process through two crucial modules: (1) Pairwise Trajectory Sampling and (2) Advantage-based Batch Shuffle.
3.1 Preliminaries
Policy gradient algorithm is a widely used RL method for MLLM, aiming to maximize the expectation of return from the environment. For a given query $q$ , $N$ independent responses $O=\{o_{1},o_{2},...,o_{N}\}$ are sampled from the old policy model $o\sim\pi_{\theta^{\prime}}(q)$. Each response receives corresponding reward $R=\{r_{1},r_{2},...,r_{N}\}$ computed via verifiable reward functions. Advantages $\hat{A}=\{\hat{A}_{1},\hat{A}_{2},...,\hat{A}_{N}\}$ are then estimated to guide policy updates. The core objective is defined as:
$\mathcal{J}(\theta)=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{N}\sim\pi_{\theta^{\prime}}(\cdot|q)}\frac{1}{\sum_{i=1}^{N}|o_{i}|}\sum_{i=1}^{N}\sum_{t=1}^{|o_{i}|}\left\{\min\left[\gamma_{t}(\theta)\hat{A}_{i},\text{clip}\left(\gamma_{t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right]\right\},$ (1)
@@ 3.1 Preliminaries
$\gamma_{t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta^{\prime}}(o_{i,t}|q,o_{i,<t})},\qquad\hat{A_{i}}=\frac{r_{i}-\text{mean}(R)}{\text{std}(R)}$ (2)
and $\epsilon$ is a clipping hyperparameter to prevent training collapse.
Despite its practicality, this static RL paradigm has notable drawbacks: advantages often concentrate near zero, yielding weak gradient signals, and the fraction of rollouts with non-trivial updates diminishes as training progresses. These issues highlight the need for a more flexible, dynamic and efficient training framework.
@@ 3.2 Problem Analysis
Our probe analysis reveals that, contrary to the ideal scenario, most rollouts exhibit advantages sharply concentrated around zero in standard RL paradigm, leading to the Advantage Collapsing phenomenon. This concentrated distribution weakens gradient signals, as only a few rollouts with high-magnitude advantages drive meaningful updates. While simply increasing the number of rollouts can partially mitigate this issue by increasing the chance of sampling valuable trajectories (Fig. 3(a)), it substantially increases computational overhead without addressing the root cause. These findings highlight the need for a dynamic mechanism that adaptively selects valuable rollouts to improve training efficiency.
Figure 3: (a) Model accuracy improves with larger rollout sizes. (b) Queries with different difficulties have varying accuracy during training, yielding rollouts with different diversity and qualities.
@@
We further observe a notable Rollout Silencing phenomenon, where the fraction of rollouts contributing non-zero gradients notably declines during training. It arises from factors such as zero advantages, gradient clipping, and excessive truncation, exposing the limits of static sampling paradigm. Tracking accuracy of queries in different difficulty reveals that simple queries converge early while difficult ones remain inaccurate (Fig. 3(b)), both failing to generate informative rollouts. Moreover, standard pipelines use each rollout only once, preventing full exploitation of valuable data. Consequently, it is crucial to design a dynamic strategy that discards ineffective rollouts while reusing informative ones.
@@ 3.3 Pairwise Trajectory Sampling
To mitigate Advantage Collapsing, we seek to select trajectories that offer stronger learning signals. Inspired by the observation that a larger rollout pool increases the probability of capturing high-advantage samples, we propose Pairwise Trajectory Sampling (PTS), a data-centric module to selectively amplify valuable learning signals. Rather than evaluating trajectories in isolation, PTS organizes candidate rollouts into structured contrastive pairs. This pairing mechanism captures both high and low advantage signals jointly, forming informative “positive-negative” pairs. Only pairs with the largest advantage contrast are then retained for training. This process ensures that limited update bandwidth is focused on trajectories that are both diverse and gradient-rich.
Given a query $q$ and a rollout size of $2N$, the rollout trajectories group is denoted as $O=\{o_{i}\}_{i=1}^{2N}$. The corresponding reward and advantage sets are $R=\{r_{i}\}_{i=1}^{2N}$ and $A=\{\hat{A}_{i}\}_{i=1}^{2N}$, respectively. To identify informative trajectory pairs, our proposed pairing mechanism follows a straightforward ‘max-min’ pairing principle by matching the trajectory with the highest advantage to that with the lowest, the second highest to the second lowest, and so on. We denote the sorted advantage values in descending order as:
@@
In this scheme, the original $2N$ rollouts are easily sorted and reorganized into $N$ pairs. The top-ranked pairs typically consist of trajectories with high-magnitude but opposite-sign advantages, forming contrastive pairs akin to ‘positive-negative’ samples. In contrast, the bottom-ranked pairs involve trajectories with advantages closer to zero.
@@
where $\alpha$ is a hyperparameter controlling the sampling ratio from the pairing set. Only the trajectories within this valid set $P_{v}$ are used for the subsequent gradient update.
By introducing a structured contrastive sampling scheme, PTS enables more effective trajectory selection from a broader exploration space without increasing the gradient computation cost. The contrastive structure not only filters out low-signal trajectories, but also sharpen the model’s policy gradient through direct comparison. PTS shifts the focus of RL fine-tuning from uniform exploration to gradient-informed selection, representing a principled step toward more efficient data usage in RL training.